Two cells each having an emf of 1.5v and an internal resistance of 2 ohms are connected in parallel. Find the current in each case when the cells are connected to a 1 ohm’s resister
1
Expert's answer
2020-11-09T06:50:45-0500

Here Two cells connected in parallel and then "1\varOmega" connected in series is analysed completely. A circuit diagram based on the question is as follows,





There are multiple ways to analyse the circuit . Here I am using Thevenins theorm.

To find the thevenins equivalent of parallel connected cells, "1\varOmega" resistor is removed and Open circuit voltage across is taken as Thevenins voltage "(V_{th})" and resistance across, shorting the voltage sources is taken as Thevenins resistance "(R_{th})" . i.e





From above circuit we have Thevenins voltage,


"V_{th} = 1.5V" (Here as there will not be any current flow through the loop . Hence the 1.5 V will show up at output port).




From the above circuit Thevenins resistance ,


"R_{th} = 2//2 = \frac{2\times2}{2+2} = 1\varOmega"


Hence we have final thevenins circuit with the load resistor "1\varOmega" as,





Hence the current through "1 \varOmega = \frac{1.5}{1+1} = 0.75A"


There fore the actual network is as follows,





As the load current is "0.75A" Current through each cell "= 0.75\/2 =0.375A"